# Stop treating ‘AGI’ as the north-star goal of AI research

### v2

@@ 1 Introduction
Recent advances in large language models (LLMs) have sparked interest in “achieving human-level ‘intelligence”’ as a “north-star goal” of the AI field (Morris et al., 2023; McCarthy et al., 1955). This goal is often referred to as “artificial general intelligence” (“AGI”) (Chollet, 2024; Tibebu, 2025). Yet rather than helping the field converge around shared goals, AGI discourse has mired it in controversies. Researchers diverge on what AGI is and assumptions about goals and risks (Summerfield, 2023; Morris et al., 2023; Blili-Hamelin et al., 2024). Researchers further contest the motivations, incentives, values, and scientific standing of claims about AGI (Gebru & Torres, 2024; Mitchell, 2024; Ahmed et al., 2024; Altmeyer et al., 2024). Finally, the building blocks of AGI as a concept—intelligence and generality—are contested in their own right (Gould, 1981; Anderson, 2002; Hernandez-Orallo & O’hEigeartiagh, 2018; Cave, 2020; Raji et al., 2021; Alexandrova & Fabian, 2022; Blili-Hamelin & Hancox-Li, 2023; Hao, 2023; Guest & Martin, 2024; Paolo et al., 2024; Mueller, 2024).
Building on prior work on the ambiguity between exploratory and confirmatory research in ML (Herrmann et al., 2024), unscientific performance claims (Altmeyer et al., 2024), SOTA-chasing (Raji et al., 2021), homogenization of research approaches (Kleinberg & Raghavan, 2021; Fishman & Hancox-Li, 2022; Bommasani et al., 2022), the values embedded in ML research (Birhane et al., 2022b), and more, our account identifies key obstacles to productive goal setting in AI research—traps [^1] Our terminology parallels Selbst et al. (2019) on fairness.. We argue that the AGI narrative makes it more difficult to overcome each obstacle. To avoid these traps, we posit that communities should stop treating AGI as a north-star goal of AI research.
@@ 2.1 Illusion of Consensus
From popular discourse to research papers to corporate marketing materials, the vast majority of references to AGI fall into this trap when they uncritically cite claims about so-called AGI. For examples of uncritical media claims, see Grossman (2023); IBM (2023); see Altmeyer et al. (2024) for examples of overhyped research. [^2] Some researchers who advocate for AGI as a goal have avoided the Illusion of Consensus trap. Morris et al. (2023) explicitly call for investigating disagreements about goals, predictions, and risks that underpin prominent accounts of AGI. Summerfield (2023) summarizes the issue: “AI researchers hope to discover how to build AGI. The problem is that nobody really knows exactly what an AGI would look like.” Mueller (2024), in turn, calls AGI “a meaningless concept, an emperor with no clothes.” Blili-Hamelin et al. (2024) identify multiple types of disagreements among attempts to define AGI or human-level AI. The contested nature of AGI as a goal becomes even more acute in critiques of AGI concepts (see, e.g., Altmeyer et al., 2024; Mueller, 2024; Van Rooij et al., 2024).
Beyond AGI, AI research is rife with topics that involve disagreement about goals, values, and concepts. For example, Mulligan et al. (2016) argue that privacy should be understood as an “essentially contested concept.” They argue that lack of agreement about the meaning and significance of privacy is not merely a matter of confusion—rather, disagreement and contestation are desirable features that enable privacy to adapt to changing technical and social contexts. Similarly, there is now widespread understanding that fairness should be understood as a contested topic, not only admitting incompatible mathematical formalizations but also incompatible values, worldviews, and theoretical assumptions (Friedler et al., 2021; Jacobs & Wallach, 2021).
@@ 2.2 Supercharging Bad Science
Underspecification of learning goals also undermines external validity—the question of whether a measurement corresponds to the real-world phenomenon it’s supposed to capture. A good example is the debate about whether “language understanding” benchmarks actually measure language understanding(Jacobs & Wallach, 2021; Liao et al., 2021).
@@
This ambiguity invites experimenter and confirmation biases, since researchers are incentivized to “pay little or no attention to competing hypotheses or explanations” or “[fail] to articulate a sufficiently strong null hypothesis,” (Altmeyer et al., 2024). Confusion between science and engineering also manifests when it is unclear if a study is pursuing scientific goals (of explanation, hypothesis confirmation, etc.) or goals of specific engineering applications (e.g., a proof-of-concept) (Hutchinson et al., 2022). This exacerbates questions about external validity: without clear and specific experimental goals, it is easier to provide post-hoc interpretations of experiments that “support” a wide variety of goals.
@@ 2.4 Goal Lottery
Researchers have studied the role of socioeconomic factors, trends, and circumstantial factors in shaping AI research. For instance, Hooker (2021) has argued that a form of hardware lottery—the greater availability of hardware with strengths in parallel processing—was key to the resurgence of deep learning in the 2010s. [^3] On similar lottery or path dependence effects, see (Dehghani et al., 2021; Fishman & Hancox-Li, 2022; Hooker, 2024; Liebowitz & Margolis, 1995; Peacock, 2009; Bauer & Gill, 2024; Rossbach, 2023). Similarly, researchers have examined the role of incentives, socioeconomic factors, and hype cycles in AI research (Narayanan & Kapoor, 2024; Wang et al., 2024; Hicks et al., 2024; Gebru & Torres, 2024; Sartori & Bocca, 2023; Delgado et al., 2023; Raji et al., 2022; Widder & Nafus, 2023). With this trap, we focus on cases where lotteries (luck) or incentives drive the adoption of unjustified goals—goals that are inadequately supported by scientific, engineering, or societal merit.
@@
The dynamics of goal lotteries are also visible in the story of the multi-decade neglect of deep learning architectures. In this case, a research agenda was sidelined for reasons that eventually proved to be misguided from an engineering, scientific, or societal perspective (e.g., due to “gatekeeping” effects against less popular research agendas; see Siler et al. 2015). Meanwhile, the AI industry went all in on the expert systems “bubble” (Haigh, 2024). Reductions in diversity within contemporary AI research can be a sign that similar mistakes are at play (§2.6). Some recent initiatives to counter homogenization (Chollet et al., 2024) rely on operationalizing AGI through benchmarks. [^4] Chollet proposes that “We will have AGI when creating [benchmarks ‘that are easy for humans, yet impossible for AI’] becomes outright impossible” (Chollet & @fchollet, 2024). In practice, they end up as yet another benchmark: incentivizing SOTA-chasing, supercharged by intense media and marketing attention (Jones, 2025). For this reason, we remain somewhat skeptical of whether approaches like ARC (Chollet, 2019) outweigh the negative consequences of news-cycle-accelerated SOTA-chasing.
@@ 2.6 Normalized Exclusion
Problem 1: Excluding communities. Many communities are left out of meaningful participation in shaping the goals of AI research (Delgado et al., 2023). Excluding communities causes serious harm, especially to minoritized communities (Pierre et al., 2021). It also undermines the utility of end products, reduces model performance, and impedes innovation (e.g., Burt, 2004). For instance, the infamous case of facial recognition engines—such as those used by Google, Apple, and Meta—mistaking Black people for gorillas (BBC News, 2015) is still occurring more than 8 years after the problem was first identified (Grant & Hill, 2023; Appelman, 2023), with downstream impacts on surveillance and law enforcement (Jones, 2020; Pour, 2023). Similarly, selective forms of inclusion in data annotation raise ethical and practical concerns about downstream effects (Wang et al., 2022; Bertelsen et al., 2024). For other examples of exclusion or inclusion of communities impacting performance, see (Buolamwini & Gebru, 2018; Young et al., 2019; Raji et al., 2020; Andrews et al., 2024; Salavati et al., 2024; Bergman et al., 2024; Weidinger et al., 2024)). Excluding communities from meaningful feedback can also undermine societal goals, such as fostering collective legitimacy through accountability to impacted communities (Birhane et al., 2022a; Young et al., 2024; Costanza-Chock, 2020; Mikesell et al., 2013; Schulz et al., 2002).
The recent prominence of AGI discourse intensifies the existing problem of community exclusion in AI research (Kelly, 2024; Frank et al., 2017). Many proponents of AGI envision a future where AI systems perform an extraordinary range of tasks for countless communities. However, research and design processes fall short of the inclusiveness demanded by this ambitious vision. For example, December 2024 reporting suggests that OpenAI and Microsoft “signed an agreement last year stating OpenAI has only achieved AGI when it develops AI systems that can generate at least \$100 billion in profits” (Maxwell Zeff, ; OpenAI, ), a stark departure from OpenAI’s public definition of AGI (OpenAI, 2018). As several authors argue, economic value is not the only type of desirable value (Morris et al., 2023; Pierson et al., 2025; Dulka, 2022; Harrigian et al., 2023; Agrawal et al., 2022). The economic definition helps guide the engineering decisions of OpenAI. But it is questionable whether an emphasis on profits will lead to the most beneficial or useful end products or to meaningful consensus about goals, especially for minoritized groups.
@@
One such major challenge is disciplinary silos, where knowledge is inadequately shared across disciplines (Ballantyne, 2019; DiPaolo, 2022; The Royal Society & Leontidis, 2024; Stirling, 2014; Amoo et al., 2020; Stokols et al., 2003). For instance, lack of knowledge sharing could be partly responsible for low attention to the distinction between explanatory and exploratory research in ML, discussed in §2.2 (Herrmann et al., 2024).
Another challenge is epistemic hierarchies—where the expertise of some disciplines is explicitly or implicitly devalued (Knorr Cetina, 1999, 2007; Fourcade et al., 2015; Simonton, 2004). This can manifest as expert groups being limited to narrow input rather than contributing to broader research design decisions (Bertelsen et al., 2024).
AI researchers’ focus on applying their work to other domains creates another major challenge. Insufficient domain knowledge might affect the functionality of AI tools—whether they operate as advertised (Raji et al., 2022). For instance, AI tools are deployed to make predictions about future outcomes that concern human individuals, from pre-trial risk prediction and predictive policing to automated employment decisions (Wang et al., 2024). Yet inadequate evidence of effectiveness often fails to prevent predictive tools from being built, marketed, and deployed (Connealy et al., 2024; Doucette et al., 2021; Cameron, 2023).
The problem of disciplinary silos becomes particularly acute in AGI-oriented research due to two factors: its claims to be creating cognates or replacements of human intelligence, and its claims to expertise in many disciplines. In the former case, claims are often made while ignoring debates in cognitive science and psychology about the nature of intelligence (Mitchell, 2024; Guest & Martin, 2024; Van Rooij et al., 2024; Summerfield, 2023). In the latter case, achieving AGI is often framed in terms of being able to “replace” domain experts in various domains—which are then often taken up uncritically by the media without input from the domain experts themselves (see, e.g., Henshall, 2024).
@@ 4.1 Alternative View
Thoughtful attempts to address shortcomings in accounts of AGI indeed exist (Adams et al., 2012; Chollet, 2019; Morris et al., 2023; Summerfield, 2023). If prior accounts of AGI are counterproductive or flawed, why not pursue new accounts that address those flaws?
As an example of the alternative view, Morris et al. (2023) arguably mitigate the Illusion of Consensus trap by disentangling the disagreements about goals, predictions, and risks that plague prior accounts of AGI. Moreover, their proposal for a practical strategy analogous to Levels of Driving Automation standards (SAE International, 2021) could be viewed as mitigating the Presuming Value-Neutrality trap. Setting society-wide standards could be done in a way that explicitly centers specific risks that the standards address (Morris et al., 2023). In this way, the values being favored manifest in the risks that are centered by the standards.
Works like Morris et al. (2023) showcase a potential response to traps. In arguing that AGI discourse aggravates multiple standing problems, we cannot rule out the possibility of efforts that mitigate these same problems while retaining AGI as a goal. Moreover, given that lack of agreement about how to define AGI is likely to persist, as Morris et al. (2023); Summerfield (2023) and many others do, it would be especially implausible for us to presume to have a complete enough view of the landscape of possible conceptions of AGI to draw definitive conclusions.
@@ Appendix A Definitions of AGI and related concepts
Gubrud (1997) “By advanced artificial general intelligence, I mean AI systems that rival or surpass the human brain in complexity and speed, that can acquire, manipulate and reason with general knowledge, and that are usable in essentially any phase of industrial or military operations where a human intelligence would otherwise be needed. Such systems may be modeled on the human brain, but they do not necessarily have to be, and they do not have to be ”conscious” or possess any other competence that is not strictly relevant to their application. What matters is that such systems can be used to replace human brains in tasks ranging from organizing and running a mine or a factory to piloting an airplane, analyzing intelligence data or planning a battle.”
@@
Legg & Hutter (2007b)Intelligence measures an agent’s ability to achieve goals in a wide range of environments.
(Fast Company, 2010)Could a computer make a cup of coffee?” Tasks the machine to go into an “average” American home, find ingredients, and make a cup of coffee. This requires embodied AI systems. Wozniak’s test has since been included in discussions of AGI. (Goertzel et al., 2012)
@@
OpenAI (2018). “OpenAI’s mission is to ensure that artificial general intelligence (AGI)—by which we mean highly autonomous systems that outperform humans at most economically valuable work—benefits all of humanity.” December 2024 reporting suggests that OpenAI and Microsoft “signed an agreement last year stating OpenAI has only achieved AGI when it develops AI systems that can generate at least \$100 billion in profits” (Maxwell Zeff, ). If true, this is a significant departure from their former definition.
Chollet et al. (2024); Chollet (2019). Defines AGI as“a system capable of efficiently acquiring new skills and solving novel problems for which it was neither explicitly designed nor trained”. In 2019, introduced an as yet (January 2025) unsolved benchmark for incentivizing progress towards AGI thus defined. Proposes that “it’s still feasible to create unsaturated, interesting benchmarks that are easy for humans, yet impossible for AI – without involving specialist knowledge. We will have AGI when creating such evals becomes outright impossible” (Chollet & @fchollet, 2024).
@@
Summerfield (2023) We consider this account to endorse a strong form of pluralism about intelligence: natural intelligence takes many different shapes across cultures and species, serving many goals and functions that are irreducibly shaped by “the internal model by which an animal understands the world”, which itself “depends on its local environment, its embodied form, its desires and goals, and its interactions with conspecifics.” The same should be expected for “strong AI” or “AGI”. However, building on Dreyfus & Dreyfus (1986), Summerfield argues we have practical reasons to constrain the forms AI takes. The goal of AI is “to help humans in their endeavours.” To that end,“if we want to build AI systems that exhibit human-like intelligence, with whom we can interact in pursuit of human-centred goals, these agents will need to think in ways that make sense to us.” 777“[W]e are building AI to make the world a better place. But if we want AI to be useful to people, it will need to share our umwelt. If we build an AI that sees the world in a radically different way to us, its behaviour and mental states will be unintelligible. Such an agent will be at best unreliable and at worst unsafe.” (Summerfield, 2023)

### v3

@@ 1 Introduction
Recent advances in large language models (LLMs) have sparked interest in “achieving human-level ‘intelligence”’ as a “north-star goal” of the AI field (Morris et al., 2023; McCarthy et al., 1955). This goal is often referred to as “artificial general intelligence” (“AGI”) (Chollet, 2024a; Tibebu, 2025). Yet rather than helping the field converge around shared goals, AGI discourse has mired it in controversies. Researchers diverge on what AGI is and assumptions about goals and risks (Summerfield, 2023; Morris et al., 2023; Blili-Hamelin et al., 2024). Researchers further contest the motivations, incentives, values, and scientific standing of claims about AGI (Gebru & Torres, 2024; Mitchell, 2024; Ahmed et al., 2024; Altmeyer et al., 2024). Finally, the building blocks of AGI as a concept—intelligence and generality—are contested in their own right (Gould, 1981; Anderson, 2002; Hernandez-Orallo & O’hEigeartiagh, 2018; Cave, 2020; Raji et al., 2021; Alexandrova & Fabian, 2022; Blili-Hamelin & Hancox-Li, 2023; Hao, 2023; Guest & Martin, 2024; Paolo et al., 2024; Mueller, 2024).
Building on prior work on the ambiguity between exploratory and confirmatory research in ML (Herrmann et al., 2024), unscientific performance claims (Altmeyer et al., 2024), SOTA-chasing (Raji et al., 2021), homogenization of research approaches (Kleinberg & Raghavan, 2021; Fishman & Hancox-Li, 2022; Bommasani et al., 2022), the values embedded in ML research (Birhane et al., 2022b), and more, our account identifies key obstacles to productive goal setting in AI research—traps. [^1] Our terminology parallels Selbst et al. (2019) on fairness. We argue that the AGI narrative makes it more difficult to overcome each obstacle. To avoid these traps, we posit that communities should stop treating AGI as a north-star goal of AI research.
@@ 2.1 Illusion of Consensus
From popular discourse to research papers to corporate marketing materials, the vast majority of references to AGI fall into this trap when they uncritically cite claims about so-called AGI. For examples of uncritical media claims, see Grossman (2023); IBM (2023); see Altmeyer et al. (2024) for examples of overhyped research. [^2] Some researchers who advocate for AGI as a goal have avoided the Illusion of Consensus trap (§2.1); e.g., Morris et al. (2023) explicitly call for investigating disagreements about goals, predictions, and risks that underpin prominent accounts of AGI. Summerfield (2023) summarizes the issue: “AI researchers hope to discover how to build AGI. The problem is that nobody really knows exactly what an AGI would look like.” Mueller (2024), in turn, calls AGI “a meaningless concept, an emperor with no clothes.” Blili-Hamelin et al. (2024) identify multiple types of disagreement among attempts to define AGI or human-level AI. The contested nature of AGI as a goal becomes even more acute in critiques of AGI concepts (see, e.g., Altmeyer et al., 2024; Mueller, 2024; Van Rooij et al., 2024).
Beyond AGI, AI research is rife with topics that involve disagreement about goals, values, and concepts. For example, Mulligan et al. (2016) argue that privacy should be understood as an “essentially contested concept.” They argue lack of agreement about the meaning and significance of privacy is not merely a matter of confusion—rather, disagreement and contestation are desirable features that enable privacy to adapt to changing technical and social contexts. Similarly, there is now widespread understanding fairness should be understood as a contested topic, not only admitting incompatible mathematical formalizations but also incompatible values, worldviews, and theoretical assumptions (Friedler et al., 2021; Jacobs & Wallach, 2021).
@@ 2.2 Supercharging Bad Science
Underspecification of learning goals also undermines external validity—the question of whether a measurement corresponds to the real-world phenomenon it’s supposed to capture. A good example is debate about whether “language understanding” benchmarks actually measure language understanding (Jacobs & Wallach, 2021; Liao et al., 2021).
@@
This ambiguity invites experimenter and confirmation biases, since researchers are incentivized to “pay little or no attention to competing hypotheses or explanations” or “[fail] to articulate a sufficiently strong null hypothesis,” (Altmeyer et al., 2024). Confusion between science and engineering also manifests when it is unclear if a study is pursuing scientific goals—of explanation, hypothesis confirmation, etc.—or goals of specific engineering applications—e.g., a proof-of-concept (Hutchinson et al., 2022). This exacerbates questions about external validity: without clear and specific experimental goals, it is easier to provide post-hoc interpretations of experiments that “support” a wide variety of goals.
@@ 2.4 Goal Lottery
Researchers have studied the role of socioeconomic factors, trends, and circumstantial factors in shaping AI research. For instance, Hooker (2021) has argued that a form of hardware lottery—the greater availability of hardware with strengths in parallel processing—was key to the resurgence of deep learning in the 2010s. [^3] On similar lottery or path dependence effects, see Dehghani et al. (2021); Fishman & Hancox-Li (2022); Hooker (2024); Liebowitz & Margolis (1995); Peacock (2009); Bauer & Gill (2024); Rossbach (2023). Similarly, researchers have examined the role of incentives, socioeconomic factors, and hype cycles in AI research (Narayanan & Kapoor, 2024; Wang et al., 2024; Hicks et al., 2024; Gebru & Torres, 2024; Sartori & Bocca, 2023; Delgado et al., 2023; Raji et al., 2022; Widder & Nafus, 2023). With this trap, we focus on cases where lotteries (luck) or incentives drive the adoption of unjustified goals—goals that are inadequately supported by scientific, engineering, or societal merit.
@@
The dynamics of goal lotteries are also visible in the story of the multi-decade neglect of deep learning architectures. In this case, a research agenda was sidelined for reasons that eventually proved to be misguided from an engineering, scientific, or societal perspective (e.g., due to “gatekeeping” effects against less popular research agendas; see Siler et al., 2015). Meanwhile, the AI industry went all in on the expert systems “bubble” (Haigh, 2024). Reductions in diversity within contemporary AI research can be a sign that similar mistakes are at play (§2.6). Some recent initiatives to counter homogenization (Chollet et al., 2024) rely on operationalizing AGI through benchmarks. [^4] Chollet proposes that “We will have AGI when creating [benchmarks ‘that are easy for humans, yet impossible for AI’] becomes outright impossible” (2024b). In practice, they end up as yet another benchmark: incentivizing SOTA-chasing, supercharged by intense media and marketing attention (Jones, 2025). For this reason, we remain somewhat skeptical of whether approaches like ARC (Chollet, 2019) outweigh the negative consequences of news-cycle-accelerated SOTA-chasing.
@@ 2.6 Normalized Exclusion
Problem 1: Excluding communities. Many communities are left out of meaningful participation in shaping the goals of AI research (Delgado et al., 2023). Excluding communities causes serious harm, especially to minoritized communities (Pierre et al., 2021); it also undermines the utility of end products, reduces model performance, oversimplifies technical challenges (Kierans et al., 2025), and impedes innovation (e.g., Burt, 2004). For instance, the infamous case of facial recognition engines—such as those used by Google, Apple, and Meta—mistaking Black people for gorillas (BBC News, 2015) is still occurring more than 8 years after the problem was first identified (Grant & Hill, 2023; Appelman, 2023), with downstream impacts on surveillance and law enforcement (Jones, 2020; Pour, 2023). Similarly, selective forms of inclusion in data annotation raise ethical and practical concerns about downstream effects (Wang et al., 2022; Bertelsen et al., 2024). Other work on exclusion or inclusion of communities impacting performance is plentiful (see Buolamwini & Gebru, 2018; Young et al., 2019; Raji et al., 2020; Andrews et al., 2024; Salavati et al., 2024; Bergman et al., 2024; Weidinger et al., 2024). Excluding communities from meaningful feedback also undermines societal goals, like fostering collective legitimacy via accountability to impacted communities (Birhane et al., 2022a; Young et al., 2024; Costanza-Chock, 2020; Mikesell et al., 2013; Schulz et al., 2002).
The recent prominence of AGI discourse intensifies the existing problem of community exclusion in AI research (Kelly, 2024; Frank et al., 2017). Many proponents of AGI envision a future where AI systems perform an extraordinary range of tasks for countless communities. However, research and design processes fall short of the inclusiveness demanded by this ambitious vision. For example, December 2024 reporting suggests that OpenAI and Microsoft “signed an agreement last year stating OpenAI has only achieved AGI when it develops AI systems that can generate at least \$100 billion in profits” (Zeff, 2025; OpenAI, 2025), a stark departure from OpenAI’s public definition of AGI (OpenAI, 2018). As several authors argue, economic value is not the only type of desirable value (Morris et al., 2023; Pierson et al., 2025; Dulka, 2022; Harrigian et al., 2023; Agrawal et al., 2022). The economic definition helps guide the engineering decisions of OpenAI. But it is questionable whether an emphasis on profits will lead to the most beneficial or useful end products or to meaningful consensus about goals, especially for minoritized groups.
@@
One major challenge is disciplinary silos, where knowledge is inadequately shared across disciplines (Ballantyne, 2019; DiPaolo, 2022; The Royal Society & Leontidis, 2024; Stirling, 2014; Amoo et al., 2020; Stokols et al., 2003). For instance, lack of knowledge sharing could be partly responsible for low attention to the distinction between explanatory and exploratory research in ML, as discussed in §2.2 (Herrmann et al., 2024).
Another challenge is epistemic hierarchies—where the expertise of some disciplines is explicitly or implicitly devalued (Knorr Cetina, 1999, 2007; Fourcade et al., 2015; Simonton, 2004; Graziul et al., 2023). This can manifest as expert groups being limited to narrow input rather than contributing to broader research design decisions (Bertelsen et al., 2024).
AI researchers’ focus on applying their work to other domains creates another major challenge. Insufficient domain knowledge might affect the functionality of AI tools—whether they operate as advertised (Raji et al., 2022). For instance, AI tools are deployed to make predictions about future individual-level outcomes, from pre-trial risk prediction and predictive policing to automated employment decisions (Wang et al., 2024). Yet inadequate evidence of effectiveness often fails to prevent predictive tools from being built, marketed, and deployed (Connealy et al., 2024; Doucette et al., 2021; Cameron, 2023).
The problem of disciplinary silos becomes particularly acute in AGI-oriented research due to two factors: its claims to be creating cognates or replacements of human intelligence, and its claims to expertise in many disciplines. In the former case, claims are often made while ignoring debates in cognitive science and psychology about the nature of intelligence (Mitchell, 2024; Guest & Martin, 2024; Van Rooij et al., 2024; Summerfield, 2023). In the latter case, achieving AGI is often framed in terms of being able to “replace” domain experts in various domains—which are then often taken up uncritically by the media without input from domain experts themselves (e.g., Henshall, 2024).
@@ 4.1 Alternative View
Thoughtful attempts to address shortcomings in accounts of AGI indeed exist (e.g., Adams et al., 2012; Chollet, 2019; Morris et al., 2023; Summerfield, 2023). If prior accounts of AGI are counterproductive or flawed, why not pursue new accounts that address those flaws?
As an example of the alternative view, Morris et al. (2023) arguably mitigate the Illusion of Consensus trap by disentangling the disagreements about goals, predictions, and risks that plague prior accounts of AGI. Moreover, their proposal for a practical strategy analogous to Levels of Driving Automation standards (SAE International, 2021) could be viewed as mitigating the Presuming Value-Neutrality trap. Setting society-wide standards could be done in a way that explicitly centers specific risks that the standards address (Morris et al., 2023). In this way, the values being favored manifest in the risks that are centered by the standards. Such works showcase a potential response to traps. In arguing that AGI discourse aggravates multiple standing problems, we cannot rule out the possibility of efforts that mitigate these same problems while retaining AGI as a goal. Moreover, given that lack of agreement about how to define AGI is likely to persist, as Morris et al. (2023), Summerfield (2023), and many others do, it would be especially implausible for us to presume to have a complete enough view of the landscape of possible conceptions of AGI to draw definitive conclusions.
@@ Appendix A Definitions of AGI and related concepts
Gubrud (1997) “By advanced artificial general intelligence, I mean AI systems that rival or surpass the human brain in complexity and speed, that can acquire, manipulate and reason with general knowledge, and that are usable in essentially any phase of industrial or military operations where a human intelligence would otherwise be needed. Such systems may be modeled on the human brain, but they do not necessarily have to be, and they do not have to be “conscious” or possess any other competence that is not strictly relevant to their application. What matters is that such systems can be used to replace human brains in tasks ranging from organizing and running a mine or a factory to piloting an airplane, analyzing intelligence data or planning a battle.”
@@
Fast Company (2010)Could a computer make a cup of coffee?” Tasks the machine to go into an “average” American home, find ingredients, and make a cup of coffee. This requires embodied AI systems. Wozniak’s test has since been included in discussions of AGI. (Goertzel et al., 2012)
@@
OpenAI (2018). “OpenAI’s mission is to ensure that artificial general intelligence (AGI)—by which we mean highly autonomous systems that outperform humans at most economically valuable work—benefits all of humanity.” December 2024 reporting suggests that OpenAI and Microsoft “signed an agreement last year stating OpenAI has only achieved AGI when it develops AI systems that can generate at least \$100 billion in profits” (Zeff, 2025). If true, this is a significant departure from their former definition.
Chollet et al. (2024); Chollet (2019). Defines AGI as“a system capable of efficiently acquiring new skills and solving novel problems for which it was neither explicitly designed nor trained”. In 2019, introduced an as yet (January 2025) unsolved benchmark for incentivizing progress towards AGI thus defined. Proposes that “it’s still feasible to create unsaturated, interesting benchmarks that are easy for humans, yet impossible for AI – without involving specialist knowledge. We will have AGI when creating such evals becomes outright impossible” (Chollet, 2024b).
@@
Summerfield (2023) We consider this account to endorse a strong form of pluralism about intelligence: natural intelligence takes many different shapes across cultures and species, serving many goals and functions that are irreducibly shaped by “the internal model by which an animal understands the world”, which itself “depends on its local environment, its embodied form, its desires and goals, and its interactions with conspecifics.” The same should be expected for “strong AI” or “AGI”. However, building on Dreyfus & Dreyfus (1986), Summerfield argues we have practical reasons to constrain the forms AI takes. The goal of AI is “to help humans in their endeavours.” To that end,“if we want to build AI systems that exhibit human-like intelligence, with whom we can interact in pursuit of human-centred goals, these agents will need to think in ways that make sense to us,”.777“[W]e are building AI to make the world a better place. But if we want AI to be useful to people, it will need to share our umwelt. If we build an AI that sees the world in a radically different way to us, its behaviour and mental states will be unintelligible. Such an agent will be at best unreliable and at worst unsafe.” (Summerfield, 2023)